OpenStudy (anonymous):

What is the Average Rate of Change from x=-4 to x=1?

OpenStudy (anonymous):

Here is the Parabola:

OpenStudy (johnweldon1993):

So we can use the average rate of change formula \[\large \frac{f(b) - f(a)}{b - a}\] Meaning: a = -4 b = 1 f(b) = what value do we get when plugging in 'b' f(a) = what value do we get when plugging in 'a'

OpenStudy (anonymous):

Is it - 3 -1 0 1

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

Thank you for giving me b and a, I think I've got it from here. Thank you so much @johnweldon1993

OpenStudy (anonymous):

Gave you best response :) Can't believe I didn't remmeber such a simple equation..

OpenStudy (johnweldon1993):

Lol dont worry about it, there's too many formulas to memorize in math but everything can be found from basic equations Such as this one can be found by manipulating the slope formula

OpenStudy (anonymous):

Just making sure the answer is 1, correct? @midhun.madhu1987 @

OpenStudy (midhun.madhu1987):

I am sorry... am not good at this...

OpenStudy (johnweldon1993):

Well at -4 we appear to be at 4 and at 1 we appear to be at -1 so \[\large \frac{-1 - 4}{1 - (-4)} = \frac{-5}{5} = -1 \]